# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

### v1

@@ D.2 Proofs for Logistic Bandits
We propose Algorithm˜5 to replace the ONS-like subroutine in the original work. We emphasize that the learning of $\mathbf{w}_{t}$ does not suffer from improper learning concerns. Because the algorithm’s final output decision is the arm $\mathbf{x}_{t}$, rather than the estimated parameter $\mathbf{w}_{t}$. Consequently, Algorithm˜5 is free from the improper-to-proper conversion and resembles LightONS.Core.
Before proving Theorem˜7, we first present two critical lemmas that facilitate migrating the original ONS-like subroutine to Algorithm˜5.
Let $\ell_{t}(\mathbf{w})=-y_{t}\log\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-(1-y_{t})\log(1-\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}))$ be the logistic loss function, then $\nabla\ell_{t}(\mathbf{w})=\left(\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-y_{t}\right)\mathbf{x}_{t}$ and $\nabla^{2}\ell_{t}(\mathbf{w})=\sigma^{\prime}(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})\mathbf{x}_{t}\mathbf{x}_{t}^{\mathsf{T}}$. For consistency, we use the notation $\mathcal{W}=\mathcal{B}(D/2)$ as in the main text, instead of $\mathcal{W}=\mathcal{B}(S)$ as in the original work.
@@
The following lemma resembles Lemma 1 of Zhang et al. [2025], showing that Algorithm˜5 implies an OMD-like regret decomposition form similar to that of original ONS-like subroutine.
@@
Proof of Lemma˜12.
Before proceeding, we note that the descend-and-project update in Algorithm˜5 is equivalent to the OMD update, as Appendix D of Zhang et al. [2025] has shown. Specifically,
@@
$\iff\mathbf{w}_{t+1}$ $=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{B}(D/2)}\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}$
@@
$\iff\mathbf{w}_{t+1}$ $=\operatorname*{arg\,min}_{\mathbf{w}\in\mathbb{R}^{d}}\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}$
@@
When $\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2$ and the Mahalanobis projection is not performed, by Eq.˜45 we have
@@
$=\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)$
which means that Eq.˜47 holds with equality. When $\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}>kD/2$ and the Mahalanobis projection is performed, by Eq.˜46, we have
$\nabla\left(\widetilde{\ell}_{t}(\mathbf{w}_{t+1})+\frac{1}{2\eta}\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)^{\mathsf{T}}(\mathbf{w}_{t+1}-\mathbf{u})\leq 0.$
@@
$=\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)$
which means that Eq.˜47 holds.
Therefore, combining both cases, we have Eq.˜47 always holds. ∎
The runtime of Algorithm˜5 is given by the following lemma.
@@
$<\left\|\left(\lambda I+\sum_{i=1}^{t}\frac{1}{\kappa}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}=\kappa^{2}\left\|\left(\kappa\lambda I+\sum_{i=1}^{t}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}$
@@
We note that $H_{t}$ and $\widetilde{H}_{t}$ admit rank-one updates similar to those in LightONS.
@@
Therefore, by reusing the proof of Lemma˜2, we obtain that the runtime of Algorithm˜5 is $\widetilde{O}(d^{2}T+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$. ∎
Algorithm 5 LightONS for [Zhang et al., 2025]
1: Initialize $H_{0}=\epsilon I$; $\mathbf{w}=\bm{0}$.
@@
5:  $\widetilde{H}_{t}=H_{t-1}+\eta\nabla^{2}\ell_{t}(\mathbf{w}_{t})$.
7:  $\mathbf{w}_{t+1}=\begin{cases}\widehat{\mathbf{w}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}]&\text{otherwise}\end{cases}$.
8:  $H_{t}=H_{t-1}+\nabla^{2}\ell_{t}(\mathbf{w}_{t+1})$.
9:end for
Based on the preceding lemmas, we are ready to prove Theorem˜7.
Proof of Theorem˜7.
The upper-confidence-bound-based pseudo-regret analysis in [Zhang et al., 2025] primarily relies on their Theorem 1, which constitutes their Lemmas 4, 5, and 6. To prove Theorem˜7, we examine how replacing the ONS-like subroutine with Algorithm˜5 affects these lemmas except the moderate expansion of the diameter from $D$ to $D^{\prime}=kD=2D$.
  • Modifications to their Lemma 4. The original Lemma 4 is supported by the original Lemma 1 and local relaxation of generalized linear models. Since the local relaxation is independent of the specific update rules of $\mathbf{w}_{t}$, we only need to verify that the original Lemma 1 in still holds when using LightONS as a replacement, as demonstrated in Lemma˜12.
  • Modifications to their Lemma 5 and 6. The original Lemma 5 and 6 depend solely on the structure of the covariance matrix, i.e., $H_{t}=\lambda I+\sum_{i=1}^{t}\sigma^{\prime}(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{w}_{i+1})\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$, and are independent of the specific update rules of $\mathbf{w}_{t}$. Therefore, plugging LightONS does not affect the original Lemma 5 and 6.
Finally, the runtime of Algorithm˜5 follows from Lemma˜13. With the overhead $O(d^{2}KT)$ for selecting arms as in [Zhang et al., 2025], the total runtime is $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$. ∎
D.3 Proofs for Memory-Efficient OXO
We propose LightONS.Sketch in Algorithm˜6, which integrates LightONS with sketching. The key difference between LightONS.Sketch and LightONS is the storage strategy. Instead of storing full matrices $A_{t}\in\mathbb{R}^{d\times d}$ and $V_{t}\in\mathbb{R}^{d\times d}$, LightONS.Sketch maintains compact sketches $S_{t}\in\mathbb{R}^{2d^{\prime}\times d}$ and $R_{t}\in\mathbb{R}^{2d^{\prime}\times 2d^{\prime}}$. Each row of $S_{t}$ stores a principal gradient component, while $R_{t}$ plays a role analogous to $V_{t}$ in Eq.˜18. The full matrix and its inverse are reconstructed as
$(A+BCD)^{-1}=A^{-1}-A^{-1}B(C^{-1}+DA^{-1}B)^{-1}DA^{-1}.$
Following [Luo et al., 2016], LightONS.Sketch uses Fast Frequent Directions [Ghashami et al., 2016] in Algorithm˜7 to update $S_{t}$ and $R_{t}$.
Before proving Theorem˜8, we first introduce three lemmas to characterize the error introduced by sketching. Specifically, Lemma˜14 bound the error for regret analysis with Lemma˜5, while Lemma˜15 bound the error for projection-count analysis with Lemma˜6. Lemma˜16 bounds the total error with the spectrum of the Hessian-related matrix.
The notation $\Delta_{t}$ in the lemmas denotes the error introduced by sketching at the $t$-th round. When SVD is performed $\Delta_{t}=\frac{2d^{\prime}}{\epsilon}\sigma_{d^{\prime}}(S_{t})^{2}$; Otherwise, $\Delta_{t}=0$. Here $\sigma_{i}(S_{t})$ is the $i$-th greatest singular value of $S_{t}$.
Lemma 14.
Proof of Lemma˜14.
Let $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$ denote the inner product induced by the Frobenius matrix norm, let $(\cdot)_{:,i}$ denote the $i$-th column of the matrix, and let $\bar{U}_{t}\bar{\Sigma}_{t}\bar{V}_{t}^{\mathsf{T}}=S_{t}$ be the full SVD with all $2d^{\prime}$ singular values. When SVD is not performed, then $\widetilde{A}_{t}=\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$ and $\Delta_{t}=0$ trivially holds. When SVD is performed,
$\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}-\widetilde{A}_{t}\right\rangle_{\operatorname{F}}=\sum_{i=1}^{2d^{\prime}}\min\left\{(\bar{\Sigma}_{t})_{i,i}^{2},(\bar{\Sigma}_{t})_{d^{\prime},d^{\prime}}^{2}\right\}\left\|(\bar{V}_{t})_{:,i}\right\|_{\widetilde{A}_{t}^{-1}}^{2}\leq\underbrace{\frac{2d^{\prime}}{\epsilon}(\bar{\Sigma}_{t})_{d^{\prime},d^{\prime}}^{2}}_{\Delta_{t}}.$
Lemma 15.
We omit proof of Lemma˜15, as it directly reuses the proof of Lemma˜14.
Lemma 16 (Theorem 1.1 and Section 3 of Ghashami et al. [2016]).
In Algorithm˜6, the total error $\Delta_{1:T}$ satisfies that, with $\lambda_{i}$ is the $i$-th greatest eigenvalue, for any $j\in[d^{\prime}]$,
Algorithm 6 LightONS.Sketch
0: preconditioner coefficient $\epsilon$, hysteresis coefficient $k$, dimension to reduce to $d^{\prime}$.
1: Initialize $\gamma^{\prime}=\frac{1}{2}\min\left\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\right\}$; $S_{0}=O_{2d^{\prime}\times d}$; $R_{0}=\frac{1}{\epsilon}I_{2d^{\prime}\times 2d^{\prime}}$; $\mathbf{x}_{1}=\mathbf{y}_{1}=\bm{0}$.
2:for $t=1,\dots,T$ do
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$.
4:  Construct $\nabla g_{t}(\mathbf{y}_{t})$ satisfying Condition˜1.
@@ D.3 Proofs for Memory-Efficient OXO
6:  $\widehat{\mathbf{y}}_{t+1}=\mathbf{y}_{t}-\frac{1}{\gamma^{\prime}}\widetilde{A}_{t}^{-1}\nabla g_{t}(\mathbf{y}_{t})$, where $\widetilde{A}_{t}^{-1}=\frac{1}{\epsilon}(I-S_{t}^{\mathsf{T}}R_{t}S_{t})$.
7:  $\mathbf{y}_{t+1}=\begin{cases}\widehat{\mathbf{y}}_{t+1}&\text{if}~\left\|\widehat{\mathbf{y}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{\widetilde{A}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{y}}_{t+1}]&\text{otherwise}\end{cases}$, where $\widetilde{A}_{t}=\epsilon I+S_{t}^{\mathsf{T}}S_{t}$.
8:  $\mathbf{x}_{t+1}=\Pi_{\mathcal{X}}[\mathbf{y}_{t+1}]$.
9:end for
Algorithm 7 Fast Frequent Directions in [Ghashami et al., 2016]
@@
With the help these lemmas, we can prove Theorem˜7 by bounding difference between LightONS and LightONS.Sketch.
Proof of Theorem˜8.
We note that Lemma˜8 holds for Algorithm˜6, as the proof of Lemma˜8 only requires $A_{t}$ to be positive-definite. Thus by Eqs.˜38 and 13 we obtain

### v2

@@ E.2 Proofs and Details for Logistic Bandits
We remark that that Algorithm˜5 does not suffer from improper learning concerns, because the algorithm’s decision is the arm $\mathbf{x}_{t}$, rather than the estimated parameter $\mathbf{w}_{t}$. Consequently, Algorithm˜5 is free from the improper-to-proper conversion and resembles LightONS.Core.
@@
Before proving Theorem˜6, we first present two critical lemmas that facilitate migrating the original ONS-based subroutine to LightONS-based Algorithm˜5. The following lemma resembles our Lemma˜11 and Lemma 1 of Zhang et al. [2025], showing that Algorithm˜5 implies an OMD-like regret decomposition form similar to that of original ONS-based subroutine.
@@
Proof of Lemma˜13.
We note that the descend-and-project update in Algorithm˜5 is equivalent to the OMD update, as Appendix D of Zhang et al. [2025] has shown. Specifically,
@@
$\iff\mathbf{w}_{t+1}$ $=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{B}(D/2)}\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2},$
@@
$\iff\mathbf{w}_{t+1}$ $=\operatorname*{arg\,min}_{\mathbf{w}\in\mathbb{R}^{d}}\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}.$
@@
When $\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2$ and the Mahalanobis projection is not triggered, by Eq.˜31 we have
@@
$=\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right),$
which means that Eq.˜33 holds with equality.
When $\left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}>kD/2$ and the Mahalanobis projection is triggered, by Eq.˜32, we have
$\nabla_{\mathbf{w}=\mathbf{w}_{t+1}}\left(\widetilde{\ell}_{t}(\mathbf{w})+\frac{1}{2\eta}\left\|\mathbf{w}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right)^{\mathsf{T}}(\mathbf{w}_{t+1}-\mathbf{u})\leq 0.$
@@
$=\frac{1}{2\eta}\left(\left\|\mathbf{w}_{t}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{u}\right\|_{H_{t-1}}^{2}-\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|_{H_{t-1}}^{2}\right),$
which means that Eq.˜33 holds.
Therefore, combining both cases, we have that Eq.˜33 always holds. ∎
The next lemma gives the total runtime of Algorithm˜5.
@@
We recall that $\ell_{t}(\mathbf{w})=-y_{t}\log\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-(1-y_{t})\log(1-\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}))$ is the logistic loss function, with $\nabla\ell_{t}(\mathbf{w})=\left(\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-y_{t}\right)\mathbf{x}_{t}$ and $\nabla^{2}\ell_{t}(\mathbf{w})=\sigma^{\prime}(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})\mathbf{x}_{t}\mathbf{x}_{t}^{\mathsf{T}}$, where $\sigma(z)=1/(1+\exp(-z))$ is the sigmoid function.
@@
$<\left\|\left(\lambda I+\sum_{i=1}^{t}\frac{1}{\kappa}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}=\kappa^{2}\left\|\left(\kappa\lambda I+\sum_{i=1}^{t}\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}\right)^{-1}\mathbf{x}_{t}\right\|_{2}^{2}.$
@@
Therefore, reusing the proof of Lemma˜2, we obtain that the runtime of Algorithm˜5 is $\widetilde{O}(d^{2}T+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$. We note that $H_{t}$ and $\widetilde{H}_{t}$ also admit rank-one updates. ∎
Based on the preceding lemmas, we are ready to prove Theorem˜6.
Proof of Theorem˜6.
The upper-confidence-bound-based pseudo-regret analysis in [Zhang et al., 2025] primarily relies on their Theorem 1, which constitutes their Lemmas 4–6. To prove Theorem˜6, we examine how replacing the ONS-based subroutine with LightONS-based Algorithm˜5 affects these lemmas, except the moderate expansion of the diameter from $D$ to $D^{\prime}=kD=2D$.
  • Modifications to their Lemma 4. Their Lemma 4 is supported by their Lemma 1 and local relaxation of generalized linear models. Lemma˜13 shows that their Lemma 1 in still holds when using Algorithm˜5 as a replacement. The local relaxation is independent of the specific update rules of $\mathbf{w}_{t}$ and remains valid.
  • Modifications to their Lemma 5 and 6. Their Lemma 5 and 6 depend solely on the structure of the covariance matrix, i.e., $H_{t}=\lambda I+\sum_{i=1}^{t}\sigma^{\prime}(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{w}_{i+1})\mathbf{x}_{i}\mathbf{x}_{i}^{\mathsf{T}}$. Therefore, Algorithm˜5 does not affect their Lemma 5 and 6.
Finally, the runtime of Algorithm˜5 follows from Lemma˜14. With the overhead $O(d^{2}KT)$ for selecting arms as in [Zhang et al., 2025], the total runtime is $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$. ∎
Comparison with OQNS.
OQNS can hardly be integrated into the method of [Zhang et al., 2025], as it is tailored to the OXO protocol and does not admit a regret decomposition analogous to Lemma˜13, which is essential for eliminating the $\kappa$ of leading terms in pseudo-regret. Moreover, OQNS lacks the flexibility to accommodate customized local norms beyond those used in standard ONS. Specifically, the local norm $\widetilde{H}_{t}$ in [Zhang et al., 2025], i.e., Eq.˜30, is not a simple accumulation of gradient outer products, unlike the OQNS’s local norm in Eq.˜4. In contrast, LightONS retains the structural flexibility of ONS, enabling integration into GLB with minimal modifications.
E.3 Proofs and Details for Memory-Efficient OXO
In this subsection, we seek to reduce the working memory required to achieve the optimal $O(d\log T)$ regret in OXO, albieit conditioned on geometric properties of the problem instance.
Current results.
Luo et al. [2016] propose Sketched Online Newton Step (SON), which achieves linear-in-$d$ runtime and working memory when the sketched dimension $d^{\prime}\ll d$. The value of $d^{\prime}$ typically depends on the problem’s intrinsic dimensionality, such as the number of non-zero eigenvalues in the Hessian-related matrix. The theoretical guarantees of SON are summarized in the following proposition.
Proposition 4 (Theorem 3 of Luo et al. [2016]).
Under Assumption˜2 and additional assumptions described in Eqs.˜36 and 37, SON (Algorithms 1 and 6 of Luo et al. [2016]) satisfies that,  [^4] The $\log T$ term in SON’s runtime arises from the eigendecomposition underlying sketching. Although Theorem 3 of Luo et al. [2016] ignores it, we include it for comparison with LightONS and OQNS.
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d^{\prime}}{\gamma}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right)+\frac{\gamma\epsilon D^{2}}{8}+\frac{\Delta_{1:T}}{2\gamma},$ (34a)
Runtime $\leq O\left(d^{\prime}dT\log T\right),$ (34b)
and the working memory is $O(d^{\prime}d)$. The cumulative sketching error $\Delta_{1:T}$ can be bounded as
$\Delta_{1:T}\leq\min_{j\in[d^{\prime}]}\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla f_{t}(\mathbf{x}_{t})\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\right).$ (35)
However, SON demands additional assumptions. Its domain $\mathcal{X}_{t}$ must be an intersection of two parallel half-spaces, onto which Mahalanobis projections admit closed-form with $O(d^{2})$ time. For general bounded convex domains in OXO, SON loses its computational advantage and reverts to the high computational cost of ONS. The domain restriction further imposes a stronger assumption on loss functions. The curvature parameter $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$ in Lemma˜1 collapses to zero for the unbounded intersection-of-parallel-half-spaces domain, reducing exp-concavity to convexity. Consequently, SON assumes an explicit quadratic property mirroring Lemma˜1, rather than the standard Assumption˜3. We list these additional assumptions as follows:
    $\mathcal{X}_{t}=\{\mathbf{x}\,|\,|\mathbf{w}_{t}^{\mathsf{T}}\mathbf{x}|<D/2\},\quad\text{where}\penalty 10000\ \mathbf{w}_{t}\in\mathbb{R}^{d}\penalty 10000\ \text{is known and}\penalty 10000\ \left\|\mathbf{w}_{t}\right\|=1.$ (36)
  • Additional assumption on loss functions. For any $t\in[T]$, the curvature parameter $\gamma$, the loss function $f_{t}$, the trajectory $\{\mathbf{x}_{t}\}_{t=1}^{T}$ and the comparator $\mathbf{u}\in\bigcap_{t=1}^{T}\mathcal{X}_{t}$ satisfies that
Improvements by LightONS.
We propose LightONS.Sketch in Algorithm˜6, replacing LightONS’s accesses to the Hessian-related matrix $A_{t}$ with the sketching primitives as in SON. LightONS.Sketch combines the projection efficiency of LightONS with the memory efficiency of SON. This hybrid method extends linear-in-$d$ runtime and working memory to the standard OXO setting (Assumptions˜1, 2 and 3), while retaining the SON’s regret $O(d^{\prime}\log T)$. Theoretical guarantees of the hybrid method are summarized in the following theorem.
Theorem 7 (LightONS’s improvement for memory-efficient OXO).
Under Assumptions˜1, 2 and 3,
with $\gamma_{0}=\frac{1}{2}\min\{\frac{1}{DG},\alpha\}$, LightONS.Sketch (Algorithm˜6) satisfies that, for any $\mathbf{u}\in\mathcal{X}$,
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\frac{d^{\prime}}{\gamma_{0}}\log\left(1+\frac{G^{2}}{2d^{\prime}\epsilon}T\right)+\frac{\gamma_{0}\epsilon D^{2}}{8}+\frac{\Delta_{1:T}}{2\gamma_{0}},$ (38a)
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{\prime}d\log T\right)T+d^{\omega}\sqrt{(d+\Delta_{1:T})T/\epsilon}\log T\right),$ (38b)
and the working memory is $O(d^{\prime}d)$. With the surrogate gradient $\nabla g_{t}(\mathbf{y}_{t})$ as in Lemma˜3, the cumulative sketching error $\Delta_{1:T}$ can be bounded as
$\Delta_{1:T}\leq\min_{j\in[d^{\prime}]}\frac{2d^{\prime}}{(d^{\prime}-j+1)\epsilon}\sum_{i=j}^{d}\lambda_{i}\left(\sum_{t=1}^{T}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right).$ (39)
Algorithm 6 LightONS.Sketch
0: preconditioner coefficient $\epsilon$, deferral coefficient $k$, dimension to reduce to $d^{\prime}$.
1: Initialize $\gamma^{\prime}=\frac{1}{2}\min\{\frac{1}{c_{f}c_{g}DG},\frac{4}{c_{f}c_{g}(k+1)DG},\alpha\}$, $S_{0}=O_{2d^{\prime}\times d}$, $R_{0}=\frac{1}{\epsilon}I_{2d^{\prime}\times 2d^{\prime}}$, $\mathbf{x}_{1}=\mathbf{y}_{1}=\bm{0}$.
@@ E.3 Proofs and Details for Memory-Efficient OXO
3:  Observe $\nabla f_{t}(\mathbf{x}_{t})$; and construct $\nabla g_{t}(\mathbf{y}_{t})$ satisfying Condition˜1.
@@
[…]
7:  $\mathbf{x}_{t+1}=\Pi_{\mathcal{X}}[\mathbf{y}_{t+1}]$.
8:end for
Algorithm 7 Fast Frequent Directions of Ghashami et al. [2016]
@@
The key difference between LightONS.Sketch and LightONS is the storage strategy. Instead of storing full matrices $A_{t}\in\mathbb{R}^{d\times d}$ and $V_{t}\in\mathbb{R}^{d\times d}$, LightONS.Sketch maintains compact sketches $S_{t}\in\mathbb{R}^{2d^{\prime}\times d}$ and $R_{t}\in\mathbb{R}^{2d^{\prime}\times 2d^{\prime}}$. Each row of $S_{t}$ stores a principal gradient component, while $R_{t}$ plays a role analogous to $V_{t}$ in Eq.˜16. The full matrix and its inverse are reconstructed as
$(A+BCD)^{-1}=A^{-1}-A^{-1}B(C^{-1}+DA^{-1}B)^{-1}DA^{-1}.$
Following Luo et al. [2016], LightONS.Sketch uses Fast Frequent Directions of Ghashami et al. [2016] in Algorithm˜7 to update $S_{t}$ and $R_{t}$.
Before proving Theorem˜7, we first introduce three lemmas to characterize the error introduced by sketching. We introduce the notation $\Delta_{t}$ to denote the sketching error at the $t$-th round.
$\Delta_{t}\triangleq\begin{cases}\frac{2d^{\prime}}{\epsilon}\sigma_{d^{\prime}}(S_{t})^{2}&\text{if SVD is triggered at $t$-th round}\\ 0&\text{otherwise}\end{cases},$
where $\sigma_{i}(S_{t})$ is the $i$-th greatest singular value of $S_{t}$. Specifically, Lemma˜15 bound the error for regret analysis with Lemma˜5, while Lemma˜16 bound the error for projection-count analysis with Lemma˜6. Lemma˜17 bounds the total error with the spectrum of the Hessian-related matrix.
Lemma 15.
Proof of Lemma˜15.
Let $\left\langle\cdot,\cdot\right\rangle_{\operatorname{F}}$ denote the inner product induced by the Frobenius matrix norm, let $(\cdot)_{:,i}$ denote the $i$-th column of the matrix, and let $\bar{U}_{t}\bar{\Sigma}_{t}\bar{V}_{t}^{\mathsf{T}}=S_{t}$ be the full SVD with all $2d^{\prime}$ singular values. When SVD is not triggered, then $\widetilde{A}_{t}=\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$ and $\Delta_{t}=0$ trivially holds. When SVD is triggered,
$\quad\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}-\widetilde{A}_{t}\right\rangle_{\operatorname{F}}$
$=\sum_{i=1}^{2d^{\prime}}\min\left\{(\bar{\Sigma}_{t})_{i,i}^{2},(\bar{\Sigma}_{t})_{d^{\prime},d^{\prime}}^{2}\right\}\left\|(\bar{V}_{t})_{:,i}\right\|_{\widetilde{A}_{t}^{-1}}^{2}\leq\frac{2d^{\prime}}{\epsilon}(\bar{\Sigma}_{t})_{d^{\prime},d^{\prime}}^{2}\triangleq\Delta_{t}.$
Lemma 16.
We omit proof of Lemma˜16, as it directly reuses the proof of Lemma˜15.
Lemma 17 (Theorem 1.1 and Section 3 of Ghashami et al. [2016]).
In Algorithm˜6, $\Delta_{1:T}$ satisfies that, for any $j\in[d^{\prime}]$,
Now we can prove Theorem˜6 by bounding difference between LightONS and LightONS.Sketch with the help of the preceding lemmas characterizing the sketching error.
Proof of Theorem˜7.
We note that Lemma˜11 holds for Algorithm˜6, as the proof of Lemma˜11 only requires $A_{t}$ to be positive-definite. Thus by Lemmas˜4 and 11 we obtain